OpenStudy (help!!!!):

A has 900 people, B has 500, and C has 600. If each division is reduced proportionately until there are 1500 people, how many people will remain in C?

OpenStudy (help!!!!):

@TheSmartOne @tom982

OpenStudy (anonymous):

So the total population is A+B+C=900+500+600=2000. This needs to be scaled down to 1500, which is a reduction of 500 or 25%. Reduce C by 25% to get your answer.

OpenStudy (help!!!!):

450!

OpenStudy (anonymous):

Spot on, nice job.

OpenStudy (daniel.ohearn1):

If you reduce c by 25% you will be in turn reducing a and b by greater proportions. So 450 can't be exactly right.

OpenStudy (daniel.ohearn1):

B and C are directly proportional to A such that A+(5/9)A+(2/3)A= 15 A=648

OpenStudy (daniel.ohearn1):

Now solve for C

OpenStudy (daniel.ohearn1):

A=642 not 648

OpenStudy (anonymous):

No, you'll also be reducing A and B by 25% too, the same proportion. A=0.75*900=675. B=0.75*500=375. C=0.75*600=450. New A+B+C=675+375+450=1500 as required.

OpenStudy (anonymous):

Old ratio A:B:C = 9:5:6 New ratio A:B:C = 675:375:450 = 9:5:6 (divide by 75). Hence the population is still the same proportion of A, B and C.